### v1

@@ B.2 Computational Efficiency
PEFT Method #Params [M] Time / it [s] Train Time [h] Memory [GB]
Full FT 93.7 |||\,| 104.8 1.44 14.33 52.1
Attn Tune (*) 32.5 |||\,| 43.5 1.42 19.72 51.4
AdaptFormer (**) 5.3 |||\,| 16.3 1.43 18.75 50.8
LoRA (**) 5.3 |||\,| 16.3 1.44 17.59 51.2
QLoRA (**) 5.3 |||\,| 16.3 1.50 18.34 51.1
FacT (**) 4.4 |||\,| 15.4 1.45 19.60 51.2
SSF (**) 4.3 |||\,| 15.3 1.49 16.12 53.3
Bias Tune (*) 4.2 |||\,| 15.2 1.42 8.34 50.1
LN Tune (*) 4.1 |||\,| 15.1 1.41 15.29 50.9
Freeze Encoder (*) 4.1 |||\,| 15.1 1.24 16.79 35.5
@@
Figure 8: Time until convergence (early stopping) when training $\mu$SAM and SAM on the microscopy datasets, including LIVECell, for the PEFT methods.
@@
Figure 9: Time until convergence (early stopping) when training MedicoSAM and SAM on the 6 medical datasets for the PEFT methods.
Appendix C Resource-efficient Finetuning
@@
Table 8: Allocated memory (in gigabytes) for each microscopy dataset during training, comparing full fine-tuning, LoRA, QLoRA and freezing the encoder when training from SAM and $\mu$SAM.
@@
Table 9: Training time until convergence in minutes for each microscopy dataset, comparing full fine-tuning, LoRA, QLoRA and freezing the encoder when training from SAM and $\mu$SAM. For CellSeg1 the training time amounts to approximately 30 minutes, for 300 epochs, without early stopping.
@@
Table 10: Time per iteration in seconds for each microscopy dataset, comparing full fine-tuning, LoRA, QLoRA and freezing the encoder when training from SAM and $\mu$SAM.
@@
Appendix D Ablation Study
D.1 LoRA

### v2

@@ E.2 Computational Efficiency
. PEFT Method #Params [M] Time / it [s] Train Time [h] Memory [GB] Full FT 93.7 |||\,| 104.8 1.44 14.33 52.1 Late FT 46.6 |||\,| 57.7 1.48 20.01 45.1 Late LoRA 6.42 |||\,| 17.45 1.49 17.51 43.6 Late QLoRA 6.42 |||\,| 17.45 1.53 21.31 45.8 Attn Tune (*) 32.5 |||\,| 43.5 1.42 19.72 51.4 AdaptFormer (**) 5.3 |||\,| 16.3 1.43 18.75 50.8 LoRA (**) 5.3 |||\,| 16.3 1.44 17.59 51.2 QLoRA (**) 5.3 |||\,| 16.3 1.50 18.34 51.1 FacT (**) 4.4 |||\,| 15.4 1.45 19.60 51.2 SSF (**) 4.3 |||\,| 15.3 1.49 16.12 53.3 Bias Tune (*) 4.2 |||\,| 15.2 1.42 8.34 50.1 LN Tune (*) 4.1 |||\,| 15.1 1.41 15.29 50.9 Freeze Encoder (*) 4.1 |||\,| 15.1 1.24 16.79 35.5
@@
Figure 10: Time until convergence (early stopping) when training $\mu$SAM and SAM on the microscopy datasets, including LIVECell, for the PEFT methods. The dashed line in black represents the average train time taken by each PEFT method.
@@
Figure 11: Time until convergence (early stopping) when training MedicoSAM and SAM on the 6 medical datasets for the PEFT methods. The dashed line in black represents the average train time taken by each PEFT method.
E.3 PEFT Recommendations
After comparing the improvements in segmentation quality from different PEFT methods, compared to full finetuning (App. E.1) and their computational efficiency (App. E.2), we summarize our overall findings on PEFT for biomedical segmentation with SAM:
  • PEFT methods that update the attention matrix weights, namely LoRA, Attention Tuning and their variants, lead to the most consistent improvement in segmentation quality, with improvements close to full finetuning.
  • The newly introduced “Late” PEFT methods, which unfreeze parameters or insert adaptation layers only in the late ViT layers, yield almost as good results as their counterparts (i.e. adaptation applied to all ViT layers).
  • PEFT does not lead to better segmentation quality than full finetuning, even when using only very small training datasets (see also App. F). Note that this observation is not trivial; for limited data one might expect advantages of less trainable parameters due to reduced overfitting.
  • Only late PEFT yields meaningful improvements in computational efficiency during training, due to the fact that SAM finetuning is activation-bound rather than parameter bound (see also App. A).
Given these observations, we give the following recommendations for finetuning SAM for biomedical segmentation:
  • With limited hardware resources (CPU): Use full freezing of the image encoder (Freeze Encoder); as this is the only viable option in this setting.
  • With medium hardware resources (Consumer-grade GPU): Use Late Finetuning (Late FT), which provides the best quality-efficiency trade-off. Other Late PEFT approaches (e.g. Late LoRA) do not provide a meaningful improvement in efficiency and do not yield to better results.
  • With high-end hardware resources (Server GPU): Use full finetuning, as PEFT does not provide any advantages in this setting.
While our recommendations where only tested for finetuning SAM for biomedical segmentation, we believe that they are also relevant for other tasks that involve ViT finetuning, as they will likely also be activation-bound rather than parameter-bound.
Appendix F Resource-efficient Finetuning
@@
[…]
fig:data_scaling Refer to caption
Figure 12: Data scaling experiments: SAM models are trained on n training images and evaluated on the same test set.
Table 8: Allocated memory (in gigabytes) for each microscopy dataset during training on one image, comparing full fine-tuning, LoRA, QLoRA and freezing the encoder when training from SAM and $\mu$SAM.
@@
Table 9: Training time on one training image until convergence in minutes for each microscopy dataset, comparing full fine-tuning, LoRA, QLoRA and freezing the encoder when training from SAM and $\mu$SAM. For CellSeg1 the training time amounts to approximately 30 minutes, for 300 epochs, without early stopping.
@@
Table 10: Time per iteration in seconds on one training image for each microscopy dataset, comparing full fine-tuning, LoRA, QLoRA and freezing the encoder when training from SAM and $\mu$SAM.
@@
[…]
fig:lateft_microscopy Refer to caption
Figure 13: Quantitative results of $\mu$SAM models trained with late finetuning and late LoRA across different settings on 6 different microscopy datasets.
[…]
fig:lateft_medical Refer to caption
Figure 14: Quantitative results of MedicoSAM models trained with late finetuning and late LoRA across different settings on 9 different medical datasets.
Appendix G Ablation Study
G.1 LoRA